OpenStudy (anonymous):

x^3+6x^2+7x-2 find all zeros

OpenStudy (anonymous):

1 doesn't work -1 doesn't work 2 doesn't work try -2

OpenStudy (anonymous):

once you see that \(f(-2)=0\) factor as \[x^3+6x^2+7x-2=(x+2)(something)\] the "something will be quadratic, use the quadratic formula or complete the square to find the other two zeros

OpenStudy (anonymous):

I actually found that -2 works, now they are asking me to factor (x^2+4x-1) and to use quadratic formula and I got 1+/- radical 5, but it seems to be wrong when I answer it in webassign

OpenStudy (anonymous):

you made a mistake in solving

OpenStudy (anonymous):

i can see from by eyeball that since you have a middle term of \(4x\) that it should not be \(1\pm\sqrt5\) but rather \(-2\pm\sqrt5\)

OpenStudy (anonymous):

\[-4\pm 2\sqrt{5}/2 is what I should have \right? I'm just confused from that point on \to get the answer\]

OpenStudy (anonymous):

that is what you get for using the quadratic formula in any case divide all by 2 and get the answer i wrote above

OpenStudy (anonymous):

\[\frac{-4\pm2\sqrt5}{2}=-2\pm\sqrt5\]

OpenStudy (anonymous):

ahh, I totally forgot that you can reduce everything, rather than factoring out the 2 from the numerator only, Because that would leave -2+-sqrt5/2, hence my answer -1+-sqrt5. Thanks a lot! I appreciate it.